OpenStudy (guccidothedishes):

CHECK MY WORK : A point Q on a segment with endpoints A(2, −1) and C(4, 2) partitions the segment in a 3:1 ratio. Find Q.

OpenStudy (guccidothedishes):

3:1 ratio means there's 4 equal parts. distance between 2 and 4 is 2. to find the 3/1 distance multiply by 2 3/1 * 2/1 = 6/1 2 (3/4) = 1.53 (3/4) = 2.25 (2+1.5,-1+2.25) =(3.5, 1.25)

OpenStudy (guccidothedishes):

^my answer

OpenStudy (amistre64):

A(2, −1) and C(4, 2) A-A, and C-A A'(0,0) and C'(2, 3) multply by 3/4 C'(2*3/4, 3*3/4) = C'(3/2, 9/4) Q = C'+A Q(3/2+2, 9/4-1) Q(7/2, 5/4) seems good to me

OpenStudy (guccidothedishes):

Thank you!